OpenStudy (anonymous):

simplify the expression p - 1/2 = 2/3will give medal to best answer.

OpenStudy (anonymous):

\[p - \frac{ 1 }{ 2 } = \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

p - 1/2 = 2/3 there are 2 ways to do this... (1) p - 1/2 = 2/3 (multiply the equation by the LCD, which is 6) 6p - 3 = 4 6p = 4 + 3 6p = 7 p = 7/6 (2) p - 1/2 = 2/3 (add 1/2 to both sides) p = 2/3 + 1/2 (common denominator is 6) p = 4/6 + 3/6 = 7/6 either way, you will arrive at the same answer :)

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

no, thank you very much.

OpenStudy (anonymous):

no problem :)